OpenStudy (acal21):

nsin(pi/n), does this series converge or diverge , if it converges then what is the limit??

ganeshie8 (ganeshie8):

what have you tried so far ?

OpenStudy (acal21):

well, i input it into the calculator and it goes from 0 to 3.1 and that is it , it seems that doesnt go any higher than that. i can figure out how to solve it

ganeshie8 (ganeshie8):

Notice that sin(pi/n) dances back and forth between -1 and 1

ganeshie8 (ganeshie8):

so nsin(pi/n) ranges netween -n and n

ganeshie8 (ganeshie8):

as n-> infinity, clearly the value of nsin(pi/n) wont be 0. what does that tell you about the series ?

OpenStudy (acal21):

ok so its an alternating series, but i checked the calculator and it does not alternate from negative to postive, they are all positive

ganeshie8 (ganeshie8):

its not alternating series

ganeshie8 (ganeshie8):

sin(pi/n) can take any value between -1 and 1 not just -1 and 1

ganeshie8 (ganeshie8):

find the limit and see what you get

ganeshie8 (ganeshie8):

\[\large \lim\limits_{n\to\infty} n\sin(\pi/n) = ?\]

OpenStudy (acal21):

it seems that there is no limit

ganeshie8 (ganeshie8):

\[\large \lim\limits_{n\to\infty} n\sin(\pi/n) = \pi\lim\limits_{n\to\infty} \dfrac{\sin(\pi/n)}{\pi/n} = \pi.1 = \pi \ne 0\]

ganeshie8 (ganeshie8):

so the series diverges

OpenStudy (acal21):

yes

ganeshie8 (ganeshie8):

the partial sums are positive and keep on increasing because sin(pi/n) is always positive for n > 1

OpenStudy (acal21):

i thought sin(pi/n) was inbetween 1 and -1

ganeshie8 (ganeshie8):

sin(x) ranges between -1 and 1 but lets look at sin(pi/n) for n>=1 : n=1, sin(pi/n) = ? n=2, sin(pi/n) = ? n=3, sin(pi/n) = ?

ganeshie8 (ganeshie8):

evaluate them to see why sin(pi/n) is positive for n>1

OpenStudy (acal21):

1-0, 2-1 ,3-.866